Singing Fish Fine Foods has a current annual cash dividend policy of $2.75. The price of the stock is set to yield a return of 14%.

What is the price of this stock if the dividend will be paid

a. for 10 ​years?

b. for 15 ​years?

c. for 40 years?

d. for 50 ​years?

e. for 100 years?

f. forever?

(Please show a full equation for computing one of the answers)
Answer #1

Price = Dividend × (1 – 1/(1 + r)n ) / r

a. Price = $2.75 × (1 – 1/(1.14)10 / 0.14 = $14.34

b. Price = $2.75 × (1 – 1/(1.14)15 / 0.14 = $16.89

c. Price = $2.75 × (1 – 1/(1.14)40 / 0.14 = $19.54

d. Price = $2.75 × (1 – 1/(1.14)50 / 0.14 = $19.61

e. Price = $2.75 × (1 – 1/(1.14)100 / 0.14 = $19.64

f. Price = $2.75 / 0.14 = $19.64